0.3(2r-3)=0.2r+0.9
Answer by faceoff57(82)

0.3(2r-3)=0.2r+0.9
0.6r - 0.9 = 0.2r + 0.9
0.6r - 0.2r = 0.9 + 0.9
0.4r = 1.8
r = 1.8/0.4
r = 4.5